OpenStudy (anonymous):

The triangular sail on Richard's boat needs to be replaced. Find the amount of fabric needed for a new sail that has a base of 16 meters and a height of 9 meters. The answer is 72 meters or 72 square meters?

OpenStudy (anonymous):

You are multiplying a metre by a metre thus the unit is metres squared

OpenStudy (whpalmer4):

Put the units in: \[A = (1/2)bh = (1/2)(16m)(9m) = 8m*9m = 72m^2\]